# Learning Task-Agnostic Skill Bases to Uncover Motor Primitives in Animal Behavior

### v1

@@ Abstract
Animals flexibly recombine a finite set of core motor primitives to meet diverse task demands, but existing behavior-segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions. To reflect the animal behavior generation procedure, we introduce skill-based imitation learning (SKIL) for behavior understanding, a reinforcement learning-based imitation framework that (1) infers interpretable skill sets, i.e., latent basis functions of behavior, by leveraging representation learning on transition probabilities, and (2) parameterizes policies as dynamic mixtures of these skills. We validate our approach on a simple grid world, a discrete labyrinth, and unconstrained videos of freely moving animals. Across tasks, it identifies reusable skill components, learns continuously evolving compositional policies, and generates realistic trajectories beyond the capabilities of traditional discrete models. By exploiting generative behavior modeling with compositional representations, our method offers a concise, principled account of how complex animal behaviors emerge from dynamic combinations of fundamental motor primitives.
@@ 1 Introduction
Animal behavior is composed of recurring patterns that reflect underlying motor primitives like back grooming, running, and sniffing. Behavior segmentation methods [1, 2, 3, 4, 5] seek to uncover structured patterns in behavior by dividing continuous pose trajectories into discrete units known as stereotyped behavioral syllables. These segmented units can then be used to build structured representations of movement for downstream neurobehavioral analysis.
@@
However, existing behavior segmentation methods face three major limitations. First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Second, they often abstract actions into high-level syllables, overlooking shared motion components across body parts. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior. Third, most segmentation models are either non-generative (e.g., clustering [4, 5]) or rely on restrictive generative assumptions (e.g., linear dynamics and hidden Markov model [1, 2, 3]), often leading to unrealistic synthesized behaviors.
@@
To address these limitations, we introduce a new perspective: modeling behavior as emerging from an agent following a reinforcement learning (RL) policy. We study behavioral dynamics and motor primitives by inferring the animal’s policy through an RL–based imitation learning framework. It not only enables more realistic behavior generation through RL but also allows us to discover motor primitives as reusable skill sets that drive behavior toward internal goals. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor primitives that go beyond the constraints of discrete segmentation.
Concretely, we first learn interpretable latent representations from behavior trajectories, which we refer to as skill sets. These represent low-level motion patterns that act as modular building blocks of behavior. Such skills can involve different body parts. By flexibly combining these skill sets, animals can generate a wide range of complex, high-level behaviors. For example, face grooming (forepaw-to-face movements) and body grooming (repetitive strokes along the torso) share common grooming-related skills, while also engaging distinct body parts specific to each movement. We leverage existing advances in representation learning in RL [6, 7, 8] to learn skill sets from transition probability functions. Then, we parametrize policies using these skill sets and apply imitation learning [9] to train the policies from animal behavior demonstrations.
@@
Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor skills.
Our method is referred to as SKIL, short for Skill-Based Imitation Learning. The fine-grained skill representation enables highly flexible behavioral characterization through the policy. Critically, over long behavioral trajectories, we assume an animal draws from a fixed set of core motor skills, regardless of the specific movements performed [10, 11]. The policy captures how the contribution of each skill evolves continuously over time, reflecting dynamic behavioral changes. Due to the spatial specificity of different motor components, some skills may be engaged briefly, while others persist over longer periods. Moreover, multiple skills can be active simultaneously to construct ongoing movement (Fig. 1). This flexible, compositional view of behavior cannot be achieved using traditional discrete behavioral syllables. SKIL thus offers a more nuanced understanding of how complex actions are generated from dynamic combinations of fundamental motor primitives. Such insight enables researchers to test whether specific fine-grained motor primitives depend on certain local neural circuits, offering a more precise understanding of movement control and its underlying neural mechanisms.
In the remainder of the paper, we introduce preliminaries in Sec. 2, which pave the way for our methods for animal behavior understanding in Sec. 3. We review related work in Sec. 4 and demonstrate the advantage of our methods empirically in Sec. 5. Finally, we conclude with a discussion in Sec. 6.
@@ 2 Preliminaries
Markov Decision Processes (MDP). To define skills that characterize mouse behavior, we begin by modeling the observed behavioral trajectories within the framework of MDPs. Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and the animal’s condition—for example, as defined by pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon. A policy $\pi:\mathcal{S}\times[H]\to\Delta(\mathcal{A})$ is a conditional distribution over actions given a state for each time-step. We assume that the animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior. The behavior is governed by a policy that seeks to maximize this internal reward.
@@
Imitation Learning. The goal of imitation learning(IL) is to find a policy $\pi$ that mimics the behavior of the animal demonstrations. In the offline IL setting, we cannot interact with the MDP environment to collect samples using policy $\pi$, but can only access a dataset of transitions sampled from the MDP by the expert, $\mathcal{D}=\left\{(s_{i},a_{i},s^{\prime}_{i})|(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),i=1,2,...,N\right\}$, where $\tau^{e}$ is the data distribution of state-action pairs generated by the expert which is the animal in this study.
[…]
$f^{*}=\mathop{\arg\max}_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\log\frac{f(x_{i},c_{i})}{f(x_{i},c_{i})+\sum_{j=1}^{k}f(y_{i,j},c_{i})},$ (2)
where $\{(x_{i},c_{i})\}_{i=1}^{n}$ are observed samples, $\{y_{i,j}\}_{i\in[n],j\in[k]}$ are artificially generated data sampled from a predefined noise distribution $q(y)$. Here, $k$ represents the ratio of negative to positive samples. This objective can be interpreted as a multi-class extension for distinguishing real data from noise.
3 Skill-based Imitation learning (SKIL) for Behavior Understanding
Given the MDP definition, we frame skill discovery from a control-theoretic perspective. In this view, skills are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
Definition 1 (Skill Set).
Given an arbitrary transition kernel $P(\cdot|s,a)$ in a linear (or low-rank) MDP [15, 16, 17], we can express it via a spectral decomposition:
@@ 3 Skill-based Imitation learning (SKIL) for Behavior Understanding
where $\phi:{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:{\mathcal{S}}\rightarrow\mathbb{R}^{d}$, and $q(s^{\prime})$ is some parametrized probability measure supported on the state space $\mathcal{S}$. We define the function $\phi$ as the skill set. The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
Because animal behavior (e.g., navigation, pose dynamics) follows structured, stereotyped patterns that evolve on low-dimensional manifolds, it can be effectively approximated by low-rank MDPs even in the presence of underlying nonlinearities.
Given the skill definition, substituting Eq. (3) and the linearized reward model into the Bellman equation (Eq. (1)), we get:
$\!Q(s,a)$ $=r(s,a)+\gamma\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}$
$=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (4)
where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$. Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of skill features $\phi(s,a)$, offering a convenient way to link skills to the policy. Following the maximum entropy reinforcement learning framework [18], we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy. Under this assumption, the optimal policy $\pi(a|s)$ can be shown to follow:
@@
Thereby, the skill function $\phi(s,a)$ not only spans the space of $Q$-functions but also provides a universal basis for constructing the policy $\pi(a|s)$ via Eq. (5). Since $\phi(s,a)$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward function or task. In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor skills available to the animal, while the weight vector $u$ captures task-specific modulations required to produce behavior consistent with different goals. With the skill set established as a foundational representation for both value functions and policies, we are equipped to interpret behavioral trajectories through the lens of skill alignment.
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (skill discovery), as well as $u$ (skill-based policy learning) from state-action-next-state observations. The learning procedure differs depending on the nature of the behavior data.
@@ 3.1 Discrete Version
Skill discovery. For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) [19, 20, 21] or spectral decomposition representation [7, 22] to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. The resulting skill set $\phi(s,a)$ is then used in the subsequent policy learning stage.
Skill-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. (5), using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
@@
Limitations for continuous data. While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the skill discovery step (Eq. (3)), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals [2]. For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in skill-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. (6) is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the skill representations and policy in continuous state-action spaces.

### v2

# Learning Task-Agnostic Motifs to Capture the Continuous Nature of Animal Behavior

@@ Abstract
Animals flexibly recombine a finite set of core motor motifs to meet diverse task demands, but existing behavior segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions. To better capture the continuous structure of behavior generation, we introduce motif-based continuous dynamics (MCD) discovery, a framework that (1) uncovers interpretable motif sets as latent basis functions of behavior by leveraging representations of behavioral transition structure, and (2) models behavioral dynamics as continuously evolving mixtures of these motifs. We validate MCD on a multi-task gridworld, a labyrinth navigation task, and freely moving animal behavior. Across settings, it identifies reusable motif components, captures continuous compositional dynamics, and generates realistic trajectories beyond the capabilities of traditional discrete segmentation models. By providing a generative account of how complex animal behaviors emerge from dynamic combinations of fundamental motor motifs, our approach advances the quantitative study of natural behavior.
@@ 1 Introduction
A critical direction in animal behavior research has been identifying recurring patterns, often referred to as stereotyped behavioral syllables, like back grooming, running, and sniffing, directly from large-scale behavior recordings. Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu & Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables. Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu & Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) switching-dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022). The segmented syllables can then be used to build structured representations of movement for downstream neurobehavioral study.
@@
However, existing behavior segmentation methods face four major limitations. First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Second, they often extract actions as abstract syllables that fail to capture how individual body parts contribute to different motions. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior. Third, most segmentation models are either non-generative (e.g., clustering (Hsu & Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors. Fourth and most importantly, classification/clustering-based methods ignore temporal dependencies and fail to capture motifs reflecting true pose dynamics, while dynamics-based methods impose linear or nonlinear dynamics assumptions that may not match actual animal behavior, making such assumptions potentially harmful.
@@
To address these limitations, we introduce a new perspective: modeling behavior under the reinforcement learning (RL) framework. We study behavioral dynamics and motor motifs by inferring the animal’s policy through an RL-based imitation learning (IL) framework. It not only enables more realistic behavior generation through RL but also allows us to discover reusable motor motif sets to construct a policy driven by internal rewards. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor motifs that go beyond the constraints of discrete segmentation. Note that Aldarondo et al. (2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash & Hochner, 2005). Building on this, we propose Motif-based Continuous Dynamics discovery (MCD) to parse long trajectories and uncover motifs and policies that reflect behavioral dynamics. Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2014; Ren et al., 2023; Shribak et al., 2024). These motifs correspond to low-level motion patterns serving as modular building blocks of behavior and can involve different body parts. For instance, face grooming (forepaw-to-face) and body grooming (torso strokes) share grooming motifs while engaging distinct body parts. These motifs can then be used to sufficiently represent policies that characterize complex high-level behaviors. Finally, we apply imitation learning to train motif-based policies from demonstrations. This framework leverages RL in two aspects: (1) motifs are inferred through RL-based representation learning, and (2) we use policies formed from motifs to characterize behavioral dynamics. As shown later, both aspects avoid any model assumptions while capturing motifs and policies that faithfully reflect behavioral dynamics.
@@
Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor motifs.
Another key innovation in constructing policies from motifs is that each motif’s contribution evolves continuously over time, reflecting dynamic behavioral changes. Some motifs may be brief, others prolonged, and multiple motifs can be active simultaneously to build ongoing movement (Fig. 1). This flexible, compositional view cannot be achieved with traditional discrete syllables. MCD thus provides a nuanced account of how complex actions arise from dynamic motif combinations and enables testing whether fine-grained motifs depend on specific neural circuits. Compared with prior segmentation methods, MCD offers soft segmentation that captures continuous time-varying processes rather than discrete switches.
Note that there is a rich literature in robotics on learning skills through imitation learning (Paraschos et al., 2013; Lioutikov et al., 2017; Li et al., 2017; Ajay et al., 2021; Peng et al., 2022; Kuang et al., 2025). We use RL-based imitation not to compete with existing skill learning methods, as most are not suited to the behavior segmentation task for neuroscience study in this paper. Rather, our goal is to employ the RL framework as a principled way to characterize the continuous nature of animal behavior while rendering fine-grained motor motifs, analogous to how Keypoint-MoSeq relies on an SLDS framework (Weinreb et al., 2024) and VAME relies on an autoencoder framework (Luxem et al., 2022). To summarize, we contribute to behavior understanding for neuroscience and neuroethology through the following points:
  • We introduce the first RL-based IL framework for behavior segmentation, a fundamental advance since RL naturally treats behavior as a decision-making process shaped by policies and rewards. Unlike dynamics-based methods, it explains why behaviors occur, not just how they unfold.
  • Within this framework, we propose RL-based representation learning to discover motif-based policies. The learned motifs and policies do not rely on dynamics assumptions or any model assumptions. They can faithfully characterize behavioral dynamics without the mismatch issues of prior segmentation methods.
  • Our method reveals the continuous nature of animal behavior, moving beyond the discrete segmentation assumptions of existing methods.
  • It provides a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs.
@@ 2 Preliminaries
Markov Decision Processes (MDP). To define motifs that characterize animal behaviors, we begin by modeling the observed behavioral trajectories within the framework of MDPs. Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and an animal’s condition—for example, positions of pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon. A policy $\pi:\mathcal{S}\times[H]\to\Delta(\mathcal{A})$ is a conditional distribution over actions given a state for each time-step. We assume that an animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior. The behavior is governed by a policy that seeks to maximize this internal reward.
@@
Offline Imitation Learning. We use imitation learning(IL) to find a policy $\pi$ that mimics animal behavior. In the offline IL setting, we cannot interact with the MDP environment to collect samples using policy $\pi$, but can only access a dataset of transitions sampled from the MDP by the expert, $\mathcal{D}=\left\{(s_{i},a_{i},s^{\prime}_{i})|(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),i=1,2,...,N\right\}$, where $\tau^{e}$ is the data distribution of state-action pairs generated by the expert which is the animal in this study.
3 Motif-based Continuous Dynamics (MCD) Discovery
Given the MDP definition, we frame motif discovery from a control-theoretic perspective. In this view, motifs are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
Definition 1 (Motif Set).
Given an arbitrary transition kernel $P(\cdot|s,a)$ in an MDP, we can express it via a spectral decomposition:
@@ 3 Motif-based Continuous Dynamics (MCD) Discovery
where $\phi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:\mathcal{S}\rightarrow\mathbb{R}^{d}$, and $q\in\Delta(\mathcal{S})$ is a parametrized probability distribution over the state space. We define the function $\phi$ as the motif set. The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2023; Zhang et al., 2022; Shribak et al., 2024). We adopt this approach here to define motifs given the transition kernel and define rewards accordingly. Since spectral decomposition derives latents directly from the transition kernel without model assumptions, motif learning is thus independent of model assumptions and faithfully reflects the motifs present in the behavior data.
Given the motif definition, substituting Eq. 2 and the linearized reward model into the Bellman equation (Eq. 1), we get:
$Q(s,a)=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (3)
where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$. Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of motif features $\phi(s,a)$, offering a convenient way to link motifs to the policy. Following the maximum entropy reinforcement learning framework (Haarnoja et al., 2018), we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy. Under this assumption, the optimal max-entropy policy $\pi(a|s)$ can be shown to follow:
@@
Proposition 1.
The policy in Eq. 4 is not based on any model assumption but emerges naturally as the max-entropy policy based on spectral decomposition of the transition kernel. Furthermore, the learned motifs $\phi$ can represent any max-entropy policy through an appropriate choice of $u$.
The reason we define $\phi(s,a)$ as motifs is that $\phi$ provides the linear basis for the environment transition, policies, and rewards. Policies characterize behavioral dynamics by describing action tendencies conditioned on state (how behaviors evolve). Combined with the environment dynamics $P(s^{\prime}|s,a)$, it induces the transition distribution $P(s^{\prime}|s)=\sum_{a}P(s^{\prime}|s,a)\,\pi(a|s)$, which governs the evolution of behavior trajectories, as is often directly modeled in dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024). Rewards, in turn, reflect the underlying driving factors of these trajectories (why behaviors evolve). Moreover, because $\phi$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward or task. In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor motifs available to animals, while the weight vector $u$ captures task-specific modulations required to produce behavior aligned with different goals. Thus, we can interpret behavioral trajectories through the lens of motifs $\phi$.
From Def. 1 and Prop. 1, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state. Thus, our method is assumption-free while capturing dynamics, unlike classification/clustering methods (no dynamics) or dynamics-based methods (restrictive assumptions).
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning that characterizes the continuous behavioral dynamics) from demonstrations. The learning procedure differs depending on the nature of the behavior data (i.e. discrete or continuous).
@@ 3.1 Discrete Version
Motif discovery. For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) (Golub & Reinsch, 1971; Golub & Van Loan, 2013; Trefethen & Bau, 2022) or spectral decomposition representation (Ren et al., 2023; HaoChen et al., 2021) to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. The resulting motif set $\phi(s,a)$ is then used in the subsequent policy learning stage.
Motif-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. 4, using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$: